Please help me solve logx-log(x=5)=1
Answer by lwsshak3(11628)

Please help me solve logx-log(x=5)=1
**
I will assume (x=5) is meant to be (x-5)
logx-log(x-5)=1
place under single log
log(x/(x-5)=1
convert to exponential form: (base(10) raised to log of number(1)=number x/(x-5)
10^1=x/(x-5)
x=10x-50
9x=50
x=50/9